OpenStudy (anonymous):

@iGreen Someone please help me understand this so I can do it by myself in the future! I have to figure out the solution for y = 2x – 1 and y = x + 3.

OpenStudy (butterflydreamer):

solve it by simultaneous equations :) Can you see how we're trying to find a solution for " y = 2x - 1" AND " y = x + 3" ? What they mean by "find a solution" to both of those equations is that we're basically trying to solve for x and y . Now, looking at the two equations provided, can you see that both are "y = ..... " ? This means that " 2x - 1 = x + 3 " (since there are TWO equations for y) From here onwards, you just want to solve for x. Once you find your value for x, you want to SUB your x-value back into either of the two equations " y = 2x - 1" or " y = x + 3" to find you y-value. :) In the end, you should have a value for x and a value for y.

OpenStudy (anonymous):

Sort of understand what you're getting at, but I'm still confused on like how to solve it.

OpenStudy (butterflydreamer):

okay :) maybe if i draw it out, it'll be easier for you to see.